Barnyard and Barley's circus is in town, and Lorena is going with her father. When Lorena and her father arrive at the performance, they can choose from six entrances to the ground floorof the main arena. There are two stairways and one elevator that go to the balcony. How many different paths can Lorena and her father take from outside the arena to the balcony?
Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52776)

.

    6*2*1 = 12.    ANSWER

Solved.

--------------------

I wrote it so shortly, because it is more than obviously / evidently.


///////////////////


After finding several different interpretations by other tutors,
my advise to the author of this post and the problem's composer
is to throw his composition to the closest garbage bin and do not show it to people around.



Answer by greenestamps(13198)

The wording of the problem allows at least three different interpretations, leading to three different answers.

By far the most likely interpretation is that there are 6 ways to get to the main floor (6 entrances) and then 3 ways to get to the balcony -- 2 staircases or an elevator. The answer in that case is 6*3 = 18.

ANSWER (very likely the intended one): 18

The other tutor uses a far less likely interpretation -- that after the 6 choices for the entrance to the main floor getting to the balcony requires using one of two staircases AND THEN an elevator. With that interpretation, the number of paths is 6*2*1 = 12.

And yet a third, even less likely interpretation that is possible with the way the problem is worded is that, after the 6 choices of an entrance to the main floor, reaching the balcony requires using 2 staircases one after the other and then an elevator. With that unlikely interpretation, there is only 1 way to get from the main floor to the balcony, so the number of paths is only 6.